OpenStudy (anonymous):

Find dy/dx by implicit differentiation x^2+3y^2-xy=5

OpenStudy (anonymous):

for these problems you apply log on both sides then solve

OpenStudy (anonymous):

\[2x+6y \frac{ dy }{dx }-\left( x \frac{ dy }{ dx }+y*1 \right)=0\] \[\left( 6y-x \right)\frac{ dy }{ dx }=y-2x,\frac{ dy }{dx }=\frac{ y-2x }{ 6y-x }\]

OpenStudy (anonymous):

where did you get y*1 from?

OpenStudy (anonymous):

\[\frac{ d }{dx }\left( uv \right)=uv'+vu'\]

OpenStudy (anonymous):

ahhh ok